Two sinusoidal waves travel in the same direction with the same amplitude, wavelength, and speed. Their resultant wave function is given by: y(x,t) = Ares sin(kx-ωt+π/4). If Ares = 8 cm, then the amplitude, A, of each of the original sinusoidal waves producing this resultant wave is:

1
Expert's answer
2021-12-13T16:55:15-0500

If two identical waves are traveling in the same direction, with the same frequency, wavelength and amplitude; BUT differ in phase the waves add together.   

"y = y_1 + y_2" where

"y = A sin (kx - ωt) \space \space and \space \space y_2 = A sin (kx - ωt + φ)"

"y = A sin(kx - \u03c9t) + A sin(kx - \u03c9t + \u03c6)"

Apply trig identity: "sin a + sin b = 2 cos((a-b)\/2) sin((a+b)\/2)"

"A sin ( a ) + A sin ( b ) = 2A cos((a-b)\/2) sin((a+b)\/2)"

"y = 2A cos (\u03c6 \/2) sin (kx - \u03c9t + \u03c6\/2)"

The resultant sinusoidal wave has the same frequency and wavelength as the original waves, but the amplitude has changed: 

"Amplitude \space equals \space 2A cos (φ /2)" with a phase angle of φ/2

here "φ = \pi /4"

Ares = "2A cos (φ /2) = 2Acos(\pi/8)=8 cm \to A = 4.32cm"